A hunter in a valley is trying to shoot a deer on a hill. The distance of the deer along his line of sight is 10(181)½ meters and the height of the hill is 90 meters. His gun has a muzzle velocity of 100 m/sec. Minimum how many meters above the deer should he aim his rifle in order to hit it? [g=10 m/s2]

Grade:11

3 Answers

Vikas TU
7 years ago
Apply second law of eqn. in y-  direcn. .i.e.
90 = 100*t – 0.5*g*t^2
solving we get t=1 and t= 19 s but t= 1 is perceived.
In this 1 s he would miss the deer.
So to hit it, he needs extra => 0.5*g*1^1 = 5 meter above the deer.
Pushkar prabhat
6 years ago
Height of hill=90m,. u=100m/s ,g=-10m/s2 s=ut+1/2gt2 _eq(1). here we find `t` it is =1or 19 s then we find minimum meter above form eq(1) s=5m
Vibekjyoti Sahoo
2 years ago
Apply second law of eqn. in y-  direcn. .i.e.
90 = 100*t – 0.5*g*t^2
solving we get t=1 and t= 19 s but t= 1 is perceived.
In this 1 s he would miss the deer.
So to hit it, he needs extra => 0.5*g*1^1 = 5 meter above the deer